### v3

# The Lehmer complex of a Bruhat interval

Abstract
We introduce Lehmer codes for several finite Coxeter groups, including all classical Weyl groups. These codes correspond to immersions of product of chains in the Bruhat order, which allow to associate to each lower Bruhat interval of these groups a multicomplex whose $f$-polynomial is the Poincaré polynomial of the interval. Via a general construction, we prove that these polynomials are $h$-polynomials of vertex–decomposable simplicial complexes, which we call Lehmer complexes. Moreover we provide a classification, in terms of unimodal permutations, of Poincaré polynomials of smooth Schubert varieties in flag manifolds.
Quest’è una ragna
dove il tuo cuor
casca, si lagna,
s’impiglia e muor.
Arrigo Boito, Otello
1 Introduction
In this paper we introduce the Lehmer complex of a lower Bruhat interval in a finite Coxeter system. This is a shellable simplicial complex whose $h$-polynomial is the rank–generating function of the interval. For Weyl groups, Björner and Ekedahl in [6] studied these polynomials by exploiting the geometry of the underlying Schubert varieties. We develop purely combinatorial tools to deal with the shape of lower Bruhat intervals, which in principle could be applied to any finite Coxeter group.
Our definition of Lehmer complexes relies on the notion of Lehmer code for a finite Coxeter group and on a general construction associating to a multicomplex a shellable simplicial complex whose $h$-vector is the $f$-vector of the multicomplex. Let $(W,S)$ be a finite Coxeter system and $e_{1},\ldots,e_{n}$ its exponents. We define a Lehmer code for $(W,S)$ as a bijection
$L:W\rightarrow[e_{1}+1]\times\ldots\times[e_{n}+1]$
such that $L^{-1}$ is a poset morphism between the direct product of chains $[e_{1}+1]\times\ldots\times[e_{n}+1]$ and the Bruhat order $(W,\leqslant)$, see Definition 4.2. A Lehmer code realizes any lower Bruhat interval as an order ideal of $[e_{1}+1]\times\ldots\times[e_{n}+1]$. Since order ideals of $\operatorname{\mathbb{N}}^{k}$ are precisely multicomplexes, this allows to consider lower Bruhat intervals as multicomplexes.
For a finite multicomplex $J$, Björner, Frankl and Stanley in [7] introduced a shellable simplicial complex $M(J)$, which we call $M$-complex, such that the $h$-polynomial of $M(J)$ equals the $f$-polynomial of $J$, see our Definition 3.3. Given a Lehmer code $L$ for $(W,S)$ and $w\in W$, the Lehmer complex $L_{w}(W,S)$ is the $M$-complex of the multicomplex $\{L(v):v\leqslant w\}$, see Definition 4.4. The $h$-polynomial of $L_{w}(W,S)$ equals the rank–generating function of the interval $[e,w]:=\{v\in W:v\leqslant w\}$, where $e$ is the identity of the group; this is stated in Theorem 4.5. In Definition 4.1 we introduce a Lehmer complex for any finite Coxeter group; in fact, to realize such complex for the whole group we do not need a Lehmer code. With respect to the length statistic, this is the analog of the Coxeter complex, whose $h$-polynomial is the generating function of the descent statistic.
In Theorem 6.4 we provide the following formula for the rank–generating function $h_{w}(q)$ of a lower Bruhat interval $[e,w]$ in terms of $q$-analogs depending on the maxima of the associated multicomplex:
$\sum_{v\leqslant w}q^{\ell(v)}=\sum\limits_{\varnothing\neq X\subseteq\max\{L(v):v\leqslant w\}}(-1)^{|X|+1}\prod\limits_{i=1}^{|S|}[(\wedge_{x\in X}\,x)_{i}+1]_{q}.$
By this result, we believe to be an important task to compute the maxima of the order ideal $\{L(v):v\leqslant w\}$ directly from the element $w$, without a complete knowledge of the interval.
The previous results hold whenever a Lehmer code for a finite Coxeter system is available. The Lehmer code for a symmetric group $S_{n}$ is a well–known bijection between the group and $[2]\times\ldots\times[n]$ described in terms of inversions and it is involved in the combinatorics of Schubert polynomials and their generalizations, see e.g. the paper of Fulton [12]. This notion is compatible with the Bruhat order, in the meaning of Definition 4.2 and this motivates our nomenclature. For groups of type $B_{n}$ and $D_{n}$, Lehmer codes were foreseen, althought not explicitly, by Stanley in [23, Section 7]. In Section 5 we realize explicit Lehmer codes for types $A_{n}$, $B_{n}$, $D_{n}$ and $H_{3}$ by exploiting the Coxeter structure of the groups of these types. For these classes of Coxeter systems, by the fact that $M$-sequences coincide with $f$-vectors of multicomplexes, in Proposition 4.3 we recover part of [6, Theorem E], extending it to type $H_{3}$. The existence of Lehmer codes for other finite irreducible Coxeter systems remains an open problem.
In Section 3 we study $M$-complexes, collecting the tools we need to prove the previous results. In Corollary 3.4, we prove that every linear extension of the order ideal $J$ provides a shelling of $M_{d}(J)$, recovering the fact that the rank–generating function of the order ideal $J$ coincides with the $h$-polynomial of $M_{d}(J)$. Theorem 3.7 provides a refinement of the well–known equivalence between finite $M$-sequences and $h$-vectors of shellable simplicial complexes, see [25, Theorem 3.3]. In particular
[…]
As consequence we have that $f$-vectors of simplicial complexes are $h$-vectors of completely balanced vertex–decomposable simplicial balls. Along the way, we obtain that
[…]
Since flag $M$-complexes are completely balanced and vertex–decomposable, this equality provides a further evidence of a conjecture due to Costantinescu and Varbaro [10, Conjecture 1.4].
The notion of Lehmer code $L$ for a Coxeter system $(W,S)$ leads us to consider a distinguished set of elements of $W$, which in Definition 6.1 we call $L$-principal elements. Principal elements for our Lehmer codes in type $A_{n}$ and $B_{n}$ were already considered with the name of Kempf elements. They arise in relation to desingularizations of Schubert varieties in Grassmannians [18] and to toric degenerations of Schubert varieties in flag varieties [16]. In Definition 6.6, we introduce $L$-unimodal elements. Proposition 7.10 states that the unimodal elements of the Lehmer code $L_{A_{n}}$ of Section 5.2 are exactly unimodal permutations. In Theorem 7.11 we prove that unimodal permutations classify Poincaré polynomials of smooth Schubert varieties in flag manifolds, namely in $S_{n}$ we have
$\left\{\sum_{v\leqslant w}q^{\ell(v)}:\mbox{$w$ is smooth}\right\}=\left\{\sum_{v\leqslant w}q^{\ell(v)}:\mbox{$w$ is unimodal}\right\}.$
Proposition 7.12 provides an alternative classification in terms of partitions into different parts. Similarly, the palindromic rank–generating functions of lower Bruhat intervals in type $H_{3}$ are classified by unimodal elements, specifically the ones of the Lehmer code $L_{H_{3}}$ defined in Section 5.5, see Example 6.8.
2 Notation and preliminaries
In this section we fix notation and recall some definitions useful for the rest of the paper. We drop the number 00 from the set of natural numbers, i.e. :=$\operatorname{\mathbb{N}}:=\{1,2,3,\ldots\}$. For $n\in\mathbb{N}$, define $[n]:=\left\{1,2,\ldots,n\right\}$, $[n]_{0}:=[n]\cup\{0\}$ and $[0]_{0}:=\{0\}$. For a set $X$, we denote by $|X|$ its cardinality, by $\mathcal{P}(X)$ its power set and by $X^{n}$ its $n$-th power under Cartesian product. If $x\in X^{n}$, we denote by $x_{i}$ the projection of $x$ on the $i$-th factor. The $q$-analog of $n$ is a polynomial defined by $[n]_{q}:=\sum\limits_{i=0}^{n-1}q^{i}$.
Posets. Let $(P,\leqslant)$ a finite poset. Given $x,y\in P$ such that $x\leqslant y$, define the interval $[x,y]:=\{z\in P:x\leqslant z\leqslant y\}$. A linear extension of $P$ is a total order $x_{1},\ldots,x_{r}$ of the elements of $P$ such that $x_{i}\leqslant x_{j}$ implies $i\leqslant j$, for all $i,j\in[r]$. Given two posets $(P,\leqslant_{P})$ and $(Q,\leqslant_{Q})$, a function $f:P\rightarrow Q$ is a poset morphism if $x\leqslant_{P}y$ implies $f(x)\leqslant_{Q}f(y)$. The poset $(P,\leqslant)$ is a lattice if the induced subposets $\{z\in P:z\leqslant x,z\leqslant y\}$ and $\{z\in P:z\geqslant x,z\geqslant y\}$ have a unique minimum $x\wedge y$ and a unique maximum $x\vee y$, respectively. Recall that the Cartesian product of posets is a poset under the componentwise order. A set $J\subseteq P$ is an order ideal of $P$ if $x\in J$ and $y\leqslant x$ implies $y\in J$. We say that $x\leqslant y$ is a covering relation $x\vartriangleleft y$ if $|[x,y]|=2$. A chain $x_{0}<x_{1}<\ldots<x_{n}$ of rank $n$ in $P$ is saturated if $x_{0}\vartriangleleft x_{1}\vartriangleleft\ldots\vartriangleleft x_{n}$. A saturated chain $x_{0}<x_{1}<\ldots<x_{n}$ of maximal rank is called maximal. If all maximal chains have the same rank $n$, the poset $P$ is said to be of rank $n$. If $P$ has maximum and minimum we say that $P$ is graded. The rank function $\rho:P\rightarrow\operatorname{\mathbb{N}}$ is defined by setting $\rho(x)$ equal to the rank of the order ideal $\{y\in P:y\leqslant x\}$, for every $x\in P$. The rank–generating function or Poincaré polynomial of $P$ is the polynomial $\sum_{x\in P}q^{\rho(x)}$.
Simplicial complexes and multicomplexes. Let $V$ be a finite set. A simplicial complex on a set of vertices $V$ is an order ideal $X\neq\varnothing$ of $(\mathcal{P}(V),\subseteq)$, whose elements are called faces. The maximal elements of $X$ are called facets. The dimension of $X$ is $\mathrm{max}\{|F|-1:F\in X\}$. A simplicial complex is pure of dimension $d$ if all its facets have the same cardinality $d+1$.
The $f$-vector $(f_{0},\ldots,f_{d},f_{d+1})$ of a $d$-dimensional simplicial complex $X$ is defined by
$f_{i}:=|\{F\in X:|F|=i\}|,$
for every $10\leqslant i\leqslant d+$ 1. The $h$-polynomial of a $d$-dimensional simplicial complex $X$ is defined by the equality
$\sum\limits_{i=0}^{d+1}h_{i}x^{d+1-i}=\sum\limits_{i=0}^{d+1}f_{i}(x-1)^{d+1-i}.$
The rank–generating function $\sum_{F\in X}q^{|F|}$ is the $f$-polynomial of $X$. The $h$-vector of $X$ is $(h_{0},h_{1},\ldots,h_{d+1})$.
Given a vertex $v$ of $X$, recall that the link of $v$ is the simplicial complex $\mathrm{link}_{X}(v):=\{F\setminus\{v\}:F\in X,v\in F\}$ and the deletion of $v$ is the simplicial complex $\mathrm{del}_{X}(v):=\{F\in X:v\notin F\}.$
A pure $d$-dimensional simplicial complex is
  • thin if every face of cardinality $d$ is contained in exaclty two facets. It is subthin if it is not thin and every face of cardinality $d$ is contained in at most two facets.
  • shellable if there exists a shelling order, i.e. a linear order $F_{1},\ldots,F_{r}$ on its facets such that for every $1\leqslant i<j\leqslant r$ there exist $v\in F_{j}\setminus F_{i}$ and $h<j$ such that $|F_{h}\cap F_{j}|=d$. The restriction of a facet $F_{i}$ is defined by $\mathcal{R}(F_{i}):=\{v\in F_{i}:F\setminus\{v\}\in\bigcup_{j=1}^{i-1}\mathcal{P}(F_{j})\}.$ Recall that for a shellable complex the $h$-vector satisfies $h_{j}=|\{i\in[r]:|\mathcal{R}(F_{i})|=j\}|$, for all $10\leqslant j\leqslant d+$ 1. In particular the $h$-vector of a shellable complex is non–negative.
  • vertex–decomposable if it is a simplex (possibly empty) or if there exists a vertex $v\in X$ satisfying the following properties:
    1. 1.
      $\mathrm{del}_{X}(v)$ is pure, i.e. $v$ is a shedding vertex;
    2. 2.
      $\mathrm{link}_{X}(v)$ is vertex–decomposable;
    3. 3.
      $\mathrm{del}_{X}(v)$ is vertex–decomposable.
  • balanced of type $a\in(\operatorname{\mathbb{N}}\cup\{0\})^{k}$, where $k\in\operatorname{\mathbb{N}}$ and $\sum_{i=1}^{k}a_{i}=d+1$, if there exists a partition $V=S_{1}\cup\ldots\cup S_{k}$ such that for every facet $F$ of $X$ we have $|F\cap S_{i}|=a_{i}$, for every $i\in[k]$. Notice that our definition is slightly different with respect to the definition given in [25, Chapter 3], because we allow the case $a_{i}=0$. It is completely balanced if it is balanced of type $(1,\ldots,1)$.
Let $k\in\operatorname{\mathbb{N}}$. A multicomplex is an order ideal of $\operatorname{\mathbb{N}}^{k}$. A simplicial complex is a finite multicomplex, considered as order ideal of $[2]^{k}$. The $f$-polynomial of a multicomplex is its rank–generating function. The $f$-vector of a multicomplex is the tuple of coefficients of its $f$-polynomial and it is called $M$-sequence. A numerical characterization of $M$-sequences, which coincide with Hilbert functions of a standard graded $K$-algebra, was proved by Macaulay, see [25, Theorem 2.2]. Moreover, finite $M$-sequences coincide with $h$-vectors of pure shellable simplicial complexes, see [25, Theorem 3.3].
Finite Coxeter systems. A finite Coxeter system $(W,S)$ is a finite group $W$ generated by a set $S$ of involutions, and relations $(st)^{m(s,t)}=e$, for all $s,t\in S$, where $e$ is the identity in $W$ and $m:S\times S\rightarrow\mathbb{N}$ is the so–called Coxeter matrix. For the classification of finite Coxeter systems we refer to [5, Appendix A1].
Given $X,Y\subseteq W$ and $w\in W$, we let $XY:=\{xy:x\in X,\,y\in Y\}\subseteq W$ and $wX:=\{wx:x\in X\}\subseteq W$. Analogously we define $Xw$.
We denote by $\ell:W\rightarrow\operatorname{\mathbb{N}}\cup\{0\}$ the length function. For $X\subseteq W$, we let $X(q):=\sum_{x\in X}q^{\ell(x)}$. For any $J\subseteq S$, the subgroup generated by $J$ is denoted by $W_{J}$ and we set
${}^{J}W:=\left\{w\in W:\ell(sw)>\ell(w),\,\forall\,s\in J\right\}.$
Any element $w\in W$ factorizes uniquely as $w=w_{J}({{}^{J}w})$, where ${{}^{J}w}\in{{}^{J}W}$, $w_{J}\in W_{J}$ and $\ell(w)=\ell(w_{J})+\ell({{}^{J}w})$.
One of the most important features of a Coxeter group is a natural partial order $\leqslant$ on it, called Bruhat order. It can be defined by the subword property (see [5, Chapter 2]). The Bruhat order is a refinement of the so–called weak orders, see [5, Chapter 3]. We denote by $\leqslant_{L}$ the left weak order, and by $\leqslant_{R}$ the right one. The posets $(W,\leqslant_{L})$, $(W,\leqslant_{R})$ and $(W,\leqslant)$ are graded with rank function $\ell$. Their minimum is the identity $e$. We denote by $w_{0}$ their maximum and by $w_{0}(J)$ the maximum of $W_{J}$. The exponents of a finite Coxeter system $(W,S)$ are the uniquely determined positive integers $e_{1},\ldots,e_{|S|}$ such that $W(q)=\prod_{i=1}^{|S|}[e_{i}+1]_{q}$.
The symmetric group of permutations of $n$ objects is denoted by $S_{n}$. A permutation $\sigma\in S_{n}$ can be written in one line notation as $\sigma(1)\sigma(2)\ldots\sigma(n)$. The group $S_{n}$ is a Coxeter group; its standard Coxeter presentation has generators $S=\left\{s_{1},\ldots,s_{n-1}\right\}$, where $s_{i}$ is the permutation $12\ldots(i+1)i\ldots n$, for all $i\in[n-1]$. The element of maximal length is $w_{0}=n(n-1)\ldots 21$. We say that a permutation $\sigma\in S_{n}$ avoids the pattern $\tau\in S_{k}$, with $k\leqslant n$, if there are no $1\leqslant i_{1}<i_{2}<\ldots<i_{k}\leqslant n$ such that $\sigma(i_{\tau^{-1}(1)})<\sigma(i_{\tau^{-1}(2)})<\ldots<\sigma(i_{\tau^{-1}(k)})$. Recall that the Bruhat order on $S_{n}$ represents the inclusions between Schubert varieties in the flag variety. It is well–known that smooth Schubert varieties corresponds to permutations avoiding the patterns $3412$ and $4231$.
3 Multicomplexes and M–complexes
In this section, given a finite multicomplex, we study a vertex–decomposable simplicial complex, which we call $M$-complex, whose $h$-vector equals the $f$-vector of the multicomplex. This construction was introduced by Björner, Frankl and Stanley in [7, Section 5] to prove the equivalence between finite $M$-sequences and $h$-vectors of shellable simplicial complexes. We strengthen this result proving that finite $M$-sequences coincide with $h$-vectors of vertex–decomposable simplicial balls. As we show in subsequent sections, $h$-polynomials of $M$-complexes realize the rank–generating functions of lower Bruhat intervals in several Coxeter groups.
Let $k\in\operatorname{\mathbb{N}}$ and $d:=(d_{1},\ldots,d_{k})\in\operatorname{\mathbb{N}}^{k}$. We shall define a pure simplicial complex $M_{d}$ of dimension
$\mathrm{dim}(M_{d})=\sum\limits_{i=1}^{k}d_{i}-k-1$
with number of facets equal to $\prod_{i=1}^{k}d_{i}$. For $x\in[d_{1}]\times\ldots\times[d_{k}]$ we define a set
$F_{x}:=\bigcup\limits_{i=1}^{k}\left([d_{i}]\setminus\{d_{i}+1-x_{i}\}\right)\times\{i\}.$ (1)
For example, if $k=3$, $d_{1}=d_{2}=3$ and $d_{3}=4$, we have that
$F_{(1,2,2)}=\{(1,1),(2,1),(1,2),(3,2),(1,3),(2,3),(4,3)\}.$
Definition 3.1.
Let $k\in\operatorname{\mathbb{N}}$ and $d\in\operatorname{\mathbb{N}}^{k}$. We call the simplicial complex
$M_{d}:=\left\{F_{x}:x\in[d_{1}]\times\ldots\times[d_{k}]\right\}$
the $M$-complex of degrees $d_{1},\ldots,d_{k}$.
Notice that if $d=(1,1,\ldots,1)$ then $M_{d}=\{\varnothing\}$. We call the numbers $e_{i}:=d_{i}-1$ the exponents of $M_{d}$, for all $i\in[k]$. Clearly the simplicial complex of Definition 3.1 is pure, balanced of type $(e_{1},\ldots,e_{k})$, and it has the announced dimension and number of facets.
We consider the componentwise order on the set $[d_{1}]\times\ldots\times[d_{k}]$. Its rank function is denoted by $\rho:[d_{1}]\times\ldots\times[d_{k}]\rightarrow\operatorname{\mathbb{N}}$, i.e. $x\mapsto\rho(x):=\sum_{i=1}^{k}(x_{i}-1)$. We prove now that all linear extensions of this poset are shellings for the $M$-complex $M_{d}$.
Theorem 3.2.
Let $M_{d}$ be the $M$-complex of degrees $d_{1},\ldots,d_{k}$ and $\omega:=\prod_{i=1}^{k}d_{i}$ the number of facets of $M_{d}$. If $x_{1},\ldots,x_{\omega}$ is a linear extension of $[d_{1}]\times\ldots\times[d_{k}]$, then $F_{x_{1}},\ldots,F_{x_{\omega}}$ is a shelling order for $M_{d}$.
Proof.
If $k=1$ the result is straightforward from the definition of $M_{d}$. So assume $k>1$. Let $x_{1},\ldots,x_{\omega}$ be a linear extension of $[d_{1}]\times\ldots\times[d_{k}]$; in particular $x_{1}=(1,1,\ldots,1)$. Let $i,j\in[\omega]$ be such that $1\leqslant i<j$. Then we have that either $x_{i}<x_{j}$ or there exist $a,b\in[k]$ such that $(x_{i})_{a}>(x_{j})_{a}$ and $(x_{i})_{b}<(x_{j})_{b}$. In both cases there exists $h\in[k]$ such that $(x_{i})_{h}<(x_{j})_{h}$. Notice that $(d_{h}+1-(x_{i})_{h},h)\in F_{x_{j}}\setminus F_{x_{i}}$. Let $z:=((x_{j})_{1},\ldots,(x_{j})_{h-1},(x_{i})_{h},(x_{j})_{h+1},\ldots,(x_{j})_{k})$. Since $(x_{i})_{h}<(x_{j})_{h}$, we have that $z<x_{j}$; hence there exists $r<j$ such that $z=x_{r}$. Moreover $F_{x_{j}}\setminus F_{x_{r}}=\{(d_{h}+1-(x_{i})_{h},h)\}$. This proves our statement. ∎
The simplicial complexes of the next definition are, mutatis mutandis, the ones introduced in [7, Section 5].
Definition 3.3.
Let $k\in\operatorname{\mathbb{N}}$ e $d\in\operatorname{\mathbb{N}}^{k}$. For any subset $\varnothing\neq J\subseteq[d_{1}]\times\ldots\times[d_{k}]$, the $M$-complex of $J$ is $M_{d}(J):=\{F_{x}:x\in J\}\subseteq M_{d}$, where the facet $F_{x}$ is defined in (1).
Recall that multicomplexes are precisely order ideals of $\mathbb{N}^{k}$, for some $k\in\operatorname{\mathbb{N}}$. Hence, finite multicomplexes are order ideals of $[d_{1}]\times\ldots\times[d_{k}]$, for some $d_{1},\ldots,d_{k}\in\mathbb{N}$. In the next result we prove that the $h$-polynomial of $M_{d}(J)$ is the rank–generating function of the order ideal $J$.
Corollary 3.4.
Let $\varnothing\neq J\subseteq[d_{1}]\times\ldots\times[d_{k}]$ be an order ideal. Then any linear extension of $J$ provides a shelling order of $M_{d}(J)$. Moreover, the polynomial $\sum_{x\in J}q^{\rho(x)}$ is the $h$-polynomial of the complex $M_{d}(J)$.
Proof.
Since order ideals are partial linear extensions, the first part is an immediate consequence of Theorem 3.2. Let $\omega:=|J|$ and $x_{1},\ldots,x_{\omega}$ be a linear extension of $J$; then $F_{x_{1}},\ldots,F_{x_{\omega}}$ is a shelling order of $M_{d}(J)$. The coefficient of degree $j$ of the $h$-polynomial of $M_{d}(J)$ is equal to |$|\{i\in[\omega]:|\mathcal{R}(F_{x_{i}})|=j\}|$. Let $i\in[\omega]$. The restriction of the facet $F_{x_{i}}$ is
$\mathcal{R}(F_{x_{i}})=\bigcup_{r=1}^{k}\{(d_{r}+1-a,r)\in[d_{r}]\times\{r\}:a<(x_{i})_{r}\},$
because $x_{1},\ldots,x_{\omega}$ is a linear extension. Therefore $|\mathcal{R}(F_{x_{i}})|=\rho(x_{i})$ for all $i\in[\omega]$ and this concludes the proof. ∎
Let $|J|>1$. It is easy to see that the complex $M_{d}(J)$ is thin if and only if $J=[d_{1}]\times\ldots\times[d_{k}]$; in the other cases it is subthin. From this fact, see [5, Appendix A2], we deduce the following corollary.
Corollary 3.5.
Let $J\subseteq[d_{1}]\times\ldots\times[d_{k}]$ be an order ideal such that $|J|>1$. Then the geometric realization of $M_{d}(J)$ is $PL$-homeomorphic to
  1. 1.
    the sphere ${\operatorname{\mathbb{S}}}^{\mathrm{dim}(M_{d})}$, if $J=[d_{1}]\times\ldots\times[d_{k}]$;
  2. 2.
    the ball ${\operatorname{\mathbb{B}}}^{\mathrm{dim}(M_{d})}$, otherwise.
It is well–known that a vertex–decomposable simplicial complex is shellable. In the next result, if $J\subseteq[d_{1}]\times\ldots\times[d_{k}]$ is an order ideal, we prove that the simplicial complex $M_{d}(J)$ has this stronger property.
Proposition 3.6.
Let $\varnothing\neq J\subseteq[d_{1}]\times\ldots\times[d_{k}]$ be an order ideal. Then $M_{d}(J)$ is vertex–decomposable.
Proof.
Denote by $X_{1}\star X_{2}:=\{F\cup G:F\in X_{1},G\in X_{2}\}$ the join of two simplicial complex $X_{1}$ and $X_{2}$. By [2, Proposition 2.4], a cone $X\star\{\{v\}\}$ over a vertex $v$ is vertex–decomposable if and only if $X$ is vertex–decomposable. Hence we may assume that there exists $x\in J$ such that $x_{k}=d_{k}$, i.e. $(d_{k},k)$ is a vertex of $M_{d}(J)$. In fact, if for every $x\in J$ we have $x_{k}<d_{k}$, then $M_{d}(J)$ is a cone over $(1,k)$.
If $\dim(M_{d}(J))=-1$ we have nothing to show. If $k=1$, we have that $M_{d}(J)=M_{(d_{1})}$ and it is not difficult to show that this is a vertex–decomposable simplicial complex. Assume $k>1$ and $\dim(M_{d}(J))\geqslant 0$. Define $I:=\{(x_{1},\ldots,x_{k-1})\in{\mathbb{N}}^{k-1}:x\in J\};$ it is clear that $I$ is an order ideal of $[d_{1}]\times\ldots[d_{k-1}]$. We proceed by induction on $k+\dim(M_{d}(J))$. If $d_{k}=1$, then $M_{d}(J)$ and $M_{(d_{1},\ldots,d_{k-1})}(I)$ are isomorphic as simplicial complexes and we conclude by induction. Assume now $d_{k}>1$. We have that
$\mathrm{del}_{(d_{k},k)}(M_{d}(J))=M_{(d_{1},\ldots,d_{k-1})}(I)\star\{\{(1,k),\ldots,(d_{k}-1,k)\}\}$
and in particular $(d_{k},k)$ is a shedding vertex. By iterating the cone operation and our inductive hypothesis it follows that $\mathrm{del}_{(d_{k},k)}(M_{d}(J))$ is vertex-decomposable. Finally consider the order ideal of $[d_{1}]\times\ldots\times[d_{k-1}]\times[d_{k}-1]$ defined by $I^{\prime}:=\{(x_{1},\ldots,x_{k-1},x_{k}-1):x\in I\}$. Since $\mathrm{link}_{(d_{k},k)}(M_{d}(J))$ and $M_{(d_{1},d_{2},\ldots,d_{k}-1)}(I^{\prime})$ are isomorphic as simplicial complexes, the result follows by induction. ∎
By Corollary 3.4 and the characterization of $M$-sequences as $f$-vectors of multicomplexes, non–zero finite $M$-sequences are exactly the $h$-vectors of $M$-complexes. More precisely, the following result holds.
Theorem 3.7.
Given a polynomial $P\neq 0$, the following are equivalent:
[…]
    the coefficients of $P$ form an $M$-sequence.
[…]
    $P$ is the $h$-polynomial of a vertex–decomposable simplicial ball.
Proof.
[…]
From the proof of the previous theorem, it is clear that in the case of simplicial complexes, i.e. order ideals in $[2]^{k}$, $M_{d}(J)$ is a completely balanced simplicial complex.
Corollary 3.8.
The $f$-polynomial of a simplicial complex is the $h$-polynomial of a completely balanced vertex–decomposable simplicial ball.
Recall that a simplicial complex, i.e. an order ideal $J\subseteq[2]^{k}$, is flag if either $J=[2]^{k}$ or $\rho(x)\leqslant 2$ for all $x\in\min([2]^{k}\setminus J)$. We recover [10, Proposition 4.1] thanks to Corollary 3.8 and the following proposition.
Proposition 3.9.
The simplicial complex $M_{d}(J)$ is flag if and only if $J$ is a flag simplicial complex.
Proof.
First of all we prove that if $J$ is a flag simplicial complex then $M_{d}(J)$ is flag. By our assumption we have that $J\subseteq[2]^{k}$ is an order ideal. Let $G:=\{(i_{1},j_{1}),\ldots,(i_{h},j_{h})\}$ be a non–face of $M_{d}(J)$, where $j_{1}\leqslant\ldots\leqslant j_{h}$ and $i_{t}\in[2]$ for all $t\in[h]$ and $h\geqslant 3$. For every $j\in[k]$ the set $X_{j}:=\{(1,j),(2,j)\}$ is a non–face of $M_{d}(J)$ by construction. If $X_{j}\subseteq G$ for some $j\in[k]$, then $G$ is not minimal. So assume $X_{j}\not\subseteq G$ for all $j\in[k]$; this implies $j_{1}<\ldots<j_{h}$. Let $x\in[2]^{k}$ such that $x_{t}=3-i_{t}$ if $t\in\{j_{1},\ldots,j_{h}\}$ and $x_{t}=1$ otherwise. Since $G\subseteq F_{x}$, then $F_{x}$ is not a facet of $M_{d}(J)$, i.e. $x\notin J$. Since $J$ is flag, there exists $y\leqslant x$, $y\notin J$ such that the set $A:=\{i\in[k]:y_{i}=2\}$ satisfies $1\leqslant|A|\leqslant 2$. It is clear that the set $\{(1,t):t\in A\}$ is contained in $G$ and it is not a face of $M_{d}(J)$, because otherwise there exists $z\in J$ such that $\{(1,t):t\in A\}\subseteq F_{z}$ and this implies $y\leqslant z$ and $y\in J$. It follows that every non–face of $M_{d}(J)$ of cardinality $\geqslant 3$ is not minimal, i.e. $M_{d}(J)$ is flag.
We prove now that if $M_{d}(J)$ is flag then $J$ is a flag simplicial complex. Consider $m_{i}:=\max\{x_{i}:x\in J\}$. Notice that, for all $i\in[k]$, the set $\{(1,i),\ldots,(m_{i},i)\}$ is a minimal non-face of $M_{d}(J)$ by construction. Hence, since $M_{d}(J)$ is flag, we have that $m_{i}\leqslant 2$, for all $i\in[k]$, i.e. $J$ is a simplicial complex. We need only to prove that $J$ is flag. If $\max(J)=\{(2,\ldots,2)\}$ we have nothing to prove. Then assume that $\max(J)\neq\{(2,\ldots,2)\}$. Let $x:=(x_{1},\ldots,x_{k})$ be a minimal non-face of $J$ and let $h:=|\{i\in[k]:x_{i}=2\}|$. Notice that $G:=\{(1,i_{1}),\ldots,(1,i_{h})\}$ is a non-face of $M_{d}(J)$, where $\{i\in[k]:x_{i}=2\}=\{i_{1},\ldots,i_{h}\}$ and $i_{1}<\ldots<i_{h}$. In fact, if it is a face, there exists $z\in J$ such that $G\subseteq F_{z}$ and then $z\geqslant x$, but $x\notin J$. It is also clear that $G$ is minimal, by the minimality of $x$. It follows that $h\leqslant 2$ and this concludes the proof. ∎
By Proposition 3.9 we obtain the following equality, which supports [10, Conjecture 1.4]:
[…]
4 Lehmer complexes of lower Bruhat intervals
In this section we introduce Lehmer codes for finite Coxeter groups and we study the shape of their lower Bruhat intervals whenever a Lehmer code is available.
By taking advantage of Definition 3.1, for a finite Coxeter system $(W,S)$ we introduce a shellable simplicial complex whose $h$-polynomial coincides with the rank–generating function of the group $W$ with the Bruhat order.
Definition 4.1.
The Lehmer complex $L(W,S)$ of a finite Coxeter system $(W,S)$ with exponents $e_{1},\ldots,e_{n}$ is the simplicial complex $M_{d}$, where $d:=(e_{1}+1,\ldots,e_{n}+1)$.
It is well–known that the rank–generating function of the poset $(W,\leqslant)$ is $\prod_{i=1}^{n}[e_{i}+1]_{q}$. Then, by Corollary 3.4, it equals the $h$-polynomial of the Lehmer complex of $(W,S)$.
It is worth to mention that Lehmer complexes are the analogues of Coxeter complexes with respect to the length statistic: both are shellable complexes and all the linear extensions of their underlying poset (the weak order in the case of the Coxeter complex) provide shelling orders, see [4].
The Lehmer codes of permutations provide a bijection between $S_{n}$ and $[1]\times\ldots\times[n]$. Such a bijective function is a poset morphism between $S_{n}$ with the left weak order and $[1]\times\ldots\times[n]$ ordered componentwise, see [11, Lemma 2.4]. Actually, it holds also that the inverse of this function is a poset morphism, whenever on $S_{n}$ we consider the Bruhat order. This is stated in Corollary 5.3. In this section we extend the notion of Lehmer code to finite Coxeter groups. For $n\in\{0\}\cup\operatorname{\mathbb{N}}$ we define $[n]_{0}:=\{0\}\cup[n]$. We recall that the relation $\leqslant$ on $W$ is the Bruhat order.
Definition 4.2.
Let $(W,S)$ be a finite Coxeter system with exponents $e_{1},\ldots,e_{n}$. A bijective function $L:W\rightarrow[e_{1}]_{0}\times\ldots\times[e_{n}]_{0}$ is a Lehmer code if
$L^{-1}:[e_{1}]_{0}\times\ldots\times[e_{n}]_{0}\rightarrow(W,\leqslant)$
is a poset morphism.
If $L$ is a Lehmer code for $(W,S)$ and $n:=|S|$, then necessarily
$\ell(w)=\rho(L(w))=\sum_{i=1}^{n}L(w)_{i}.$
For $w\in W$ we define
$h_{w}(q):=\sum_{v\leqslant w}q^{\ell(v)},$
the rank–generating function of the interval $[e,w]$, and
$J_{w}:=\{L(v)+(1,\ldots,1):v\leqslant w\}.$
For all finite Coxeter groups with a Lehmer code we recover in a purely combinatorial way a result due to Björner and Ekedahl, see [6, Theorem E]. They prove it for finite Weyl groups as a consequence of the geometry of their Schubert varieties.
Proposition 4.3.
Let $L$ be a Lehmer code for a finite Coxeter system $(W,S)$. Then the coefficients of $h_{w}(q)$ form an $M$-sequence, for all $w\in W$.
Proof.
Let $e_{1},\ldots,e_{n}$ be the exponents of $(W,S)$. Since the inverse of a Lehmer code is a poset morphism, the set $J_{w}$ is an order ideal of $[e_{1}+1]\times\ldots\times[e_{n}+1]$, i.e. a multicomplex. Clearly $h_{w}(q)$ equals the $f$-polynomial of $J_{w}$ and then its coefficients form an $M$-sequence. ∎
In view of the next section, the previous result applies to Coxeter systems of types $A_{n}$, $B_{n}$, $D_{n}$ and $H_{3}$.
Now we introduce a simplicial complex whose $h$-polynomial equals the rank–generating function of the lower Bruhat interval $[e,w]$.
Definition 4.4.
Let $L$ be a Lehmer code for a finite Coxeter system $(W,S)$ with exponents $e_{1},\ldots,e_{n}$; for $w\in W$ the Lehmer complex $L_{w}(W,S)$ of $[e,w]$ is the simplicial complex $M_{d}(J_{w})$, where $d:=(e_{1}+1,\ldots,e_{n}+1)$.
In view of the analogy between the complex $L(W,S)$ and the Coxeter complex of $(W,S)$, the shellable complex $L_{w}(W,S)$ is an analogue of a Coxeter cone as introduced in [26]. The next result is an immediate consequence of all the machinery developed up to this point.
Theorem 4.5.
Let $L$ be a Lehmer code for a finite Coxeter system $(W,S)$ and $w\in W$. Then the $h$-polynomial of the simplicial complex $L_{w}(W,S)$ is equal to $h_{w}(q)$.
Proof.
Since $\ell(w)=\sum_{i=1}^{n}L(w)_{i}=\rho(L(w))$, for all $w\in W$, by Corollary 3.4 the result follows. ∎
5 Lehmer codes for some finite Coxeter groups
In this section we construct several Lehmer codes for some finite irreducible Coxeter systems. In general the non–reducible case can be traced back to the irreducible one in view of the following observation. If $(W,S)\simeq(W_{1},S_{1})\times(W_{2},S_{2})$, and $L_{1}$, $L_{2}$ are Lehmer codes for $(W_{1},S_{1})$ and $(W_{2},S_{2})$, respectively, then the assignment
$w\mapsto L(w)=(L_{1}(w_{1})_{1},\ldots,L_{1}(w_{1})_{|S_{1}|},L_{2}(w_{2})_{1},\ldots,L_{2}(w_{2})_{|S_{2}|})$
defines a Lehmer code for $(W,S)$, where $w=w_{1}w_{2}\in W$, $w_{1}\in W_{1}$ and $w_{2}\in W_{2}$.
While a rank preserving bijection $W\simeq[e_{1}]_{0}\times\ldots\times[e_{n}]_{0}$ exists for all finite Coxeter systems, where $e_{1},\ldots,e_{n}$ are the exponents of $(W,S)$, the additional condition to be a poset morphism is not fullfilled by all such bijections.
Remark 5.1.
The existence of Lehmer codes for classical Weyl groups, although in different terms, was already foreseen by R. Stanley in [23, Section 7]. In [3], S. Billey has proved that, in type $A_{n}$ and $B_{n}$, any lower Bruhat interval with palindromic Poincaré polynomial has a "Lehmer code", i.e. it admits an immersed product of chains, see [3, Corollary 1.3]. We checked that this property holds also in type $H_{3}$.
Given a Lehmer code for $(W,S)$, one can construct directly several other Lehmer codes. In fact, it is clear that if $L$ is a Lehmer code for $(W,S)$ and $f$ is an automorphism of the poset $(W,\leqslant)$, then $L\circ f$ is a Lehmer code for $(W,S)$. In the dihedral system $I_{2}(m)$, the automorphism group of the Bruhat order is isomorphic to $\operatorname{\mathbb{Z}}_{2}^{m-1}$. In all other cases the automorphism group of $(W,\leqslant)$ can be described in terms of the automorphism group of the Coxeter graph of $(W,S)$ and the inversion automorphism $w\mapsto w^{-1}$, see [6, Section 2.3] and references therein. Denote by $L^{*}$ the dual of the Lehmer code $L$, which we define by setting
$L^{*}(w):=L(w^{-1}),$
for all $w\in W$. This is a Lehmer code by the previous discussion.
One of the main tool for our constructions is the machinery described in [5, Section 7.1] to factorize the Poincaré polynomial of $(W,S)$. We establish here some useful notation. Let $S=\{s_{1},\ldots,s_{n}\}$ and $w\in W$. As proved in [5, Corollary 2.4.6] (see also [5, Proposition 3.4.2]) we have a factorization
$w=w^{(1)}w^{(2)}\cdots w^{(n)},$
such that $w^{(i)}\in{{}^{\{s_{1},\ldots,s_{i-1}\}}W_{\{s_{1},\ldots,s_{i}\}}}$, for all $i\in[n]$, and $\ell(w)=\ell(w^{(1)})+\ldots+\ell(w^{(n)})$. Then the Poincaré polynomial of $(W,S)$ factorizes as
$W(q)=\prod_{i=1}^{n}{{}^{\{s_{1},\ldots,s_{i-1}\}}W_{\{s_{1},\ldots,s_{i}\}}}(q).$
For the notation on irreducible Coxter systems we follow [5, Appendix A1], sometimes with a different linear order on the generators. We also observe that for every $u,v\in W$, if $u^{(i)}\leqslant v^{(i)}$ for all $i\in[n]$ then, by the subword property of the Bruhat order, $u=u^{(1)}\cdots u^{(n)}\leqslant v^{(1)}\cdots v^{(n)}=v$. This is crucial for the construction of some Lehmer codes. In the following subsections, for sake of clarity, we adopt the same notation for the type of a Coxeter system and the group defining it.
5.1 Lehmer codes for type $I_{2}(m)$
Let $m\in\{3,4,\ldots\}$. The Coxeter system $(W,S)=I_{2}(m)$ has two generators $s,t$ which satisfy the relation $(st)^{m}=e$. The Poincaré polynomial factorizes as $W(q)=W_{\{s\}}(q)\cdot{{}^{\{s\}}W}(q)=[2]_{q}[m]_{q}$. Let
$L_{I_{2}(m)}:W\rightarrow[1]_{0}\times[m-1]_{0}$
be defined by setting $L_{I_{2}(m)}(w)=(\ell(w^{(1)}),\ell(w^{(2)}))$, for all $w\in W$. Then the function $L_{I_{2}(m)}$ is a Lehmer code for $I_{2}(m)$. It is straightforward that there are $2^{m-1}$ distinct Lehmer codes for $I_{2}(m)$ and they are all obtained by composing $L_{I_{2}(m)}$ with an automorphism of $(W,\leqslant)$.
5.2 A Lehmer code for type $A_{n}$
The Poincaré polynomial for Coxeter systems $(A_{n},S)$ of type $A_{n}$ factorizes as $A_{n}(q)=[2]_{q}[3]_{q}\cdots[n+1]_{q}$. The quotients (${{}^{\{s_{1},\ldots,s_{i-1}\}}(A_{n})_{\{s_{1},\ldots,s_{i}\}}}$ with the Bruhat order are chains, for all $i\in[n]$. Let
$L_{A_{n}}:A_{n}\rightarrow[1]_{0}\times[2]_{0}\times\ldots\times[n]_{0},$
be defined by $L_{A_{n}}(w)=(\ell(w^{(1)}),\ldots,\ell(w^{(n)}))$, for all $w\in A_{n}$. It is clear that the function $L_{A_{n}}$ is a Lehmer code for type $A_{n}$. For example, if $w=s_{2}s_{1}s_{3}s_{2}$ then $L_{A_{3}}(w)=(0,2,2)$; in fact $w^{(1)}=e$, $w^{(2)}=s_{2}s_{1}$ and $w^{(3)}=s_{3}s_{2}$.
We now describe the Lehmer code $L_{A_{n-1}}$ in terms of inversions. Consider the symmetric group $S_{n}$ as a Coxeter group of type $A_{n-1}$. Let
$L_{n}(w):=(0,L_{A_{n-1}}(w)_{1},\ldots,L_{A_{n-1}}(w)_{n-1}),$ (2)
for all $w\in S_{n}$.
Theorem 5.2.
Let $n\geqslant 2$. Then
$L_{n}(w)_{k}=|\{i\in[n]:i<k,\,w^{-1}(i)>w^{-1}(k)\}|,$
for all $k\in[n]$, $w\in S_{n}$.
Proof.
We prove the result by induction on $n$. If $n=2$ the result is trivial. Assume $n>2$. If $w(n)=n$, we have that $L_{n}(w)_{n}=0=|\{i\in[n-1]:w^{-1}(i)>w^{-1}(n)\}|$ and, by our inductive hypothesis, $L_{n}(w)_{k}=|\{i\in[k-1]:w^{-1}(i)>w^{-1}(k)\}|$, for all $k<n$.
Let $a:=w^{-1}(n)<n$, and $\widetilde{w}:=w^{(1)}\cdots w^{(n-2)}$. Then, for $k<n$, by the inductive hypothesis we have that
$L_{n}(w)_{k}$ $=$ $L_{n-1}(\widetilde{w})_{k}$
$=$ |$|\{i\in[k-1]:\widetilde{w}^{-1}(i)>\widetilde{w}^{-1}(k)\}|$
$=$ |$|\{i\in[k-1]:w^{-1}(i)>w^{-1}(k)\}|,$
since $w^{(n-1)}=s_{n-1}s_{n-2}\cdots s_{a}$ and then
$w=\tilde{w}(1)\ldots\tilde{w}(a-1)\,n\,\tilde{w}(a)\ldots\tilde{w}(n-1).$
Moreover we have that
$L_{n}(w)_{n}$ $=$ $\ell(w)-\ell(\tilde{w})$
$=$ |$|\{i\in[n-1]:w^{-1}(i)>w^{-1}(n)\}|$
and this concludes the proof. ∎
For $n\geqslant 3$, we denote by $\iota_{n}$ the non–trivial automorphism of the Coxeter graph of $A_{n-1}$. If $n=2$, $\iota_{n}$ is the identity. If $w_{0}$ is the element of maximal length in $S_{n}$, then $\iota_{n}(w)=w_{0}ww_{0}$, for all $w\in S_{n}$. In the following corollary we recover what is known as Lehmer code of a permutation, see e.g. [19, Section 11.3]. It is worth to mention that the Lehmer code of a permutation provide the leading term of the corresponding Schubert polynomial, see [12].
Corollary 5.3.
Let $n\in\operatorname{\mathbb{N}}$. Then
|$|\{j\in[n]:j>i,\,w(j)<w(i)\}|=(L_{n}\circ\iota_{n})^{*}(w)_{n-i+1},$
for all $i\in[n]$, $w\in S_{n}$.
Proof.
By Theorem 5.2 we obtain
$L_{n}(w_{0}w^{-1}w_{0})_{n-i+1}$ $=$ |$|\{j\in[n-i]:(w_{0}ww_{0})(j)>(w_{0}ww_{0})(n-i+1)\}|$
$=$ |$|\{j\in[n-i]:w(i)>w(n-j+1)\}|$
$=$ |$|\{j\in[n]:j>i,\,w(i)>w(j)\}|.$
We have that $(L_{n}\circ\iota_{n})^{*}(w)=(L_{n}\circ\iota_{n})(w^{-1})=L_{n}(w_{0}w^{-1}w_{0})$ and the result follows. ∎
By Corollary 5.3, what is known as Lehmer code satifies the condition of Definition 4.2 and then it is a Lehmer code in our meaning.
5.3 A Lehmer code for type $B_{n}$
[…]
Figure 1: Coxeter graph of $B_{n}$.
We can construct a Lehmer code for type $B_{n}$ similarly to type $A_{n}$, as we state in the next theorem. Let $(B_{n},S)$ be a Coxeter system of type $B_{n}$; we take the linear order on $S$ as in Figure 1.
Theorem 5.4.
Let $n\geqslant 2$. Then the function
$L_{B_{n}}:B_{n}\rightarrow[1]_{0}\times[3]_{0}\times\ldots\times[2n-1]_{0}$
defined by setting $L_{B_{n}}(w)=(\ell(w^{(1)}),\ldots,\ell(w^{(n)}))$, for all $w\in B_{n}$, is a Lehmer code.
Proof.
If $n\geqslant 2$, we have that the maximal quotient ${{}^{S\setminus\{s_{n}\}}B_{n}}$ is a chain of rank $12n-$ 1. In fact,
$\ell({{}^{S\setminus\{s_{n}\}}w_{0}})=\ell(w_{0})-\ell(w_{0}(S\setminus\{s_{n}\}))=n^{2}-(n-1)^{2}=2n-1$
and
$|{{}^{S\setminus\{s_{n}\}}B_{n}}|=\frac{2^{n}n!}{2^{n-1}(n-1)!}=2n.$
This implies our statement. ∎
Remark 5.5.
It is worth to mention that Lehmer codes $L_{A_{n}}$ and $L_{B_{n}}$ have been considered in [15] for the study of bases for the Bruhat order in types $A_{n}$ and $B_{n}$. Moreover, in type $B_{n}$, a signed Lehmer code is defined in [22].
5.4 A Lehmer code for type $D_{n}$
Let $n\geqslant 4$ and $(D_{n},S)$ be a Coxeter system of type $D_{n}$, whose generators are ordered as in Figure 2.
[…]
Figure 2: Coxeter graph of $D_{n}$.
We define $n$ chains in the Bruhat order of the group $D_{n}$ as follows:
$X_{1}$ ::=:= $(D_{n})_{\{s_{1}\}}$
$X_{2}$ ::=:= (${{}^{\{s_{1}\}}}(D_{n})_{\{s_{1},s_{2}\}}\cup\,s_{0}s_{2}s_{1}\{e\}$
$X_{3}$ ::=:= (${{}^{\{s_{1},s_{2}\}}}(D_{n})_{\{s_{1},s_{2},s_{3}\}}\cup\,s_{3}s_{0}s_{2}s_{1}\{e,s_{3}\}$
$\vdots$
$X_{n-1}$ ::=:= (${{}^{\{s_{1},\ldots,s_{n-2}\}}}(D_{n})_{\{s_{1},\ldots,s_{n-1}\}}$
$\cup\,\,s_{n-1}s_{n-2}\cdots s_{3}s_{0}s_{2}s_{1}\{e,s_{3},\ldots,s_{3}s_{4}\cdots s_{n-1}\}$
$Y_{n}$ ::=:= $\{e,s_{0},s_{0}s_{2},s_{0}s_{2}s_{3},\ldots,s_{0}s_{2}s_{3}\cdots s_{n-1}\}$
We need the following result to define a Lehmer code for type $D_{n}$.
Proposition 5.6.
Let $n\geqslant 4$. Then   $Y_{n-1}\cdot{{}^{S\setminus\{s_{n-1}\}}D_{n}}=X_{n-1}Y_{n}$.
Proof.
We have that $|X_{n-1}|=2(n-1)$, $|Y_{n}|=n$ and $|^{S\setminus\{s_{n-1}\}}D_{n}|=\frac{2^{n-1}n!}{2^{n-2}(n-1)!}=2n$. Hence
$|Y_{n-1}\cdot{{}^{S\setminus\{s_{n-1}\}}D_{n}}|$ $=$ $|Y_{n-1}|\cdot|{{}^{S\setminus\{s_{n-1}\}}D_{n}}|$
$=$ $(n-1)\cdot 2n=|X_{n-1}|\cdot|Y_{n}|.$
Therefore, to prove our statement, it is sufficient to prove the inclusion $Y_{n-1}\cdot{{}^{S\setminus\{s_{n-1}\}}D_{n}}\,\subseteq\,X_{n-1}Y_{n}$. First notice that
${{}^{S\setminus\{s_{n-1}\}}D_{n}}$ $=$ (${}^{S\setminus\{s_{n-1}\}}(D_{n})_{\{s_{1},\ldots,s_{n-1}\}}$
$\uplus\,\,s_{n-1}\cdots s_{2}\{s_{0},s_{1}s_{0},s_{1}s_{0}s_{2},\ldots,s_{1}s_{0}s_{2}\cdots s_{n-1}\}$
$=$ (${}^{S\setminus\{s_{n-1}\}}(D_{n})_{\{s_{1},\ldots,s_{n-1}\}}\uplus\,\,s_{n-1}\cdots s_{2}(\{s_{0}\}\cup s_{1}Y_{n}).$
Let $y\in Y_{n-1}$ and $u\in{{}^{S\setminus\{s_{n-1}\}}D_{n}}$. The previous equality implies that $u\in X_{n-1}Y_{n}$; so, if $y=e$, we have that $yu\in X_{n-1}Y_{n}$. Let $y=s_{0}$. If $u=e$ then $yu\in Y_{n-1}\subseteq Y_{n}\subseteq X_{n-1}Y_{n}$. If $u>e$ there are some cases to be considered:
  1. 1.
    $u=s_{n-1}\cdots s_{i}$, $12<i\leqslant n-$ 1: in this case $yu=uy\in X_{n-1}Y_{n}$.
  2. 2.
    $u=s_{n-1}\cdots s_{2}$: in this case $yu=s_{n-1}\cdots s_{3}\cdot s_{0}s_{2}\in X_{n-1}Y_{n}$.
  3. 3.
    $u=s_{n-1}\cdots s_{2}s_{1}$: in this case $yu=s_{n-1}\cdots s_{4}\cdot s_{3}s_{0}s_{2}s_{1}\in X_{n-1}$.
  4. 4.
    $u=s_{n-1}\cdots s_{2}s_{0}$: we have that $yu=s_{n-1}\cdots s_{3}s_{2}\cdot s_{0}s_{2}\in X_{n-1}Y_{n}$.
  5. 5.
    $u=s_{n-1}\cdots s_{2}s_{1}\cdot s_{0}s_{2}\cdots s_{i}$, $12<i<n-$ 1: we have that $yu=s_{n-1}\cdots s_{3}s_{0}s_{2}s_{1}\cdot s_{0}s_{2}\cdots s_{i}\in X_{n-1}Y_{n}$.
Now assume $\ell(y)>1$ and $u\in{{}^{\{s_{1},\ldots,s_{n-2}\}}}(D_{n})_{\{s_{1},\ldots,s_{n-1}\}}$. Then $y=s_{0}s_{2}\cdots s_{j}$ for some $22\leqslant j\leqslant n-$ 2, and $u=s_{n-1}\cdots s_{i}$, for some $i\in[n-1]$. We consider the following cases.
  1. 1.
    $12\leqslant j<i-$ 1: in this case $i>3$ and $yu=uy\in X_{n-1}Y_{n}$.
  2. 2.
    $j=i-1$: we have that $i\geqslant 3$ and $yu=s_{n-1}\cdots s_{i+1}\cdot s_{0}s_{2}\cdots s_{i}\in X_{n-1}Y_{n}$.
  3. 3.
    $i-1<j<n-1$, $i\geqslant 2$: we have that $2\leqslant i\leqslant j$ and
    $yu$ $=$ $s_{0}s_{2}\cdots s_{j}\cdot s_{n-1}\cdots s_{j+1}s_{j}\cdots s_{i}$
    $=$ $s_{n-1}\cdots s_{j+2}\cdot s_{0}s_{2}\cdots s_{i-1}(s_{i}\cdots s_{j}s_{j+1}s_{j}\cdots s_{i})$
    $=$ $s_{n-1}\cdots s_{j+2}\cdot s_{0}s_{2}\cdots s_{i-1}(s_{j+1}\cdots s_{i+1}s_{i}s_{i+1}\cdots s_{j+1})$
    $=$ $s_{n-1}\cdots s_{i+1}\cdot s_{0}s_{2}\cdots s_{i-1}s_{i}s_{i+1}\cdots s_{j+1}\in X_{n-1}Y_{n}.$
  4. 4.
    $12\leqslant j<n-$ 1, $i=1$: we have that
    $yu$ $=$ $s_{0}s_{2}\cdots s_{j}\cdot s_{n-1}\cdots s_{1}$
    $=$ $s_{n-1}\cdots s_{j+2}s_{0}(s_{2}\cdots s_{j}s_{j+1}s_{j}\cdots s_{2})s_{1}$
    $=$ $s_{n-1}\cdots s_{j+2}s_{0}(s_{j+1}\cdots s_{3}s_{2}s_{3}\cdots s_{j+1})s_{1}$
    $=$ $s_{n-1}\cdots s_{3}s_{0}s_{2}s_{1}s_{3}\cdots s_{j+1}\in X_{n-1}.$
Now let $y\in Y_{n-1}$ and $u=s_{n-1}\cdots s_{2}z$, with $z\in s_{0}\{e,s_{1},s_{1}s_{2},\ldots,s_{1}s_{2}\cdots s_{n-1}\}$. Then, by point $3$ above, we have that:
  1. 1.
    if $z=s_{0}$ then
    $yu$ $=$ $s_{n-1}\cdots s_{3}\cdot s_{0}s_{2}s_{3}\cdots s_{j+1}s_{0}$
    $=$ $s_{n-1}\cdots s_{3}s_{2}\cdot s_{0}s_{2}s_{3}\cdots s_{j+1}\in X_{n-1}Y_{n};$
  2. 2.
    if $z=s_{0}s_{1}s_{2}\cdots s_{i}$, $11\leqslant i\leqslant n-$ 1 then
    $yu$ $=$ $s_{n-1}\cdots s_{3}\cdot s_{0}s_{2}s_{3}\cdots s_{j+1}s_{0}s_{1}s_{2}\cdots s_{i}$
    $=$ $s_{n-1}\cdots s_{3}s_{0}s_{2}s_{1}s_{3}\cdots s_{j+1}\cdot s_{0}s_{2}\cdots s_{i}\in X_{n-1}Y_{n}.$
From the result of the previous proposition we deduce a factorization of the group $D_{n}$ realized by $n$ saturated chains of its Bruhat order.
Corollary 5.7.
Let $n\geqslant 4$. Then $D_{n}=X_{1}X_{2}\cdots X_{n-1}Y_{n}$.
Proof.
We claim that $D_{n}\subseteq X_{1}\cdots X_{n-2}X_{n-1}Y_{n}$, for all $n\geqslant 4$. We prove the claim by induction on $n$. Let $n=4$. We have that
$D_{4}=(D_{4})_{\{s_{1}\}}\cdot{{}^{\{s_{1}\}}(D_{4})}_{\{s_{1},s_{2}\}}\cdot{{}^{\{s_{1},s_{2}\}}(D_{4})}_{\{s_{0},s_{1},s_{2}\}}\cdot{{}^{S\setminus\{s_{3}\}}D_{4}}$
and (${{}^{\{s_{1},s_{2}\}}(D_{4})}_{\{s_{0},s_{1},s_{2}\}}=Y_{3}\cup\{s_{0}s_{2}s_{1}\}$. One can check that
(${{}^{\{s_{1}\}}}(D_{4})_{\{s_{1},s_{2}\}}\cdot s_{0}s_{2}s_{1}=\{s_{0}s_{2}s_{1},s_{2}s_{0}s_{2}s_{1},s_{2}s_{1}s_{0}s_{2}s_{1}\}\subseteq X_{2}Y_{3}.$
Therefore, by Proposition 5.6, $D_{4}\subseteq X_{1}X_{2}X_{3}Y_{4}$. Let $n>4$ and $w=w_{J}({{}^{J}w})\in D_{n}$, with $J:=S\setminus\{s_{n-1}\}$. Then $w_{J}\in D_{n-1}$ and, by our inductive hypothesis, we have that $w_{J}\in X_{1}\cdots X_{n-2}Y_{n-1}$. Hence, by Proposition 5.6, we obtain $w_{J}({{}^{J}w})\in X_{1}\cdots X_{n-2}X_{n-1}Y_{n}$ and this proves our claim. ∎
By Corollary 5.7, any element $w\in D_{n}$ is factorized uniquely as $w=x_{1}\cdots x_{n-1}y$, with $y\in Y_{n}$ and $x_{i}\in X_{i}$, for all $i\in[n-1]$. Moreover the sets $Y_{n}$ and $X_{i}$, for all $i\in[n-1]$, are saturated chains of the Bruhat order of $D_{n}$. From these facts we obtain a Lehmer code for type $D_{n}$, for all $n\geqslant 4$.
Theorem 5.8.
The function $L_{D_{n}}:D_{n}\rightarrow[1]_{0}\times[3]_{0}\times\ldots\times[2n-3]_{0}\times[n-1]_{0}$ defined by setting
$L_{D_{n}}(w)=(\ell(x_{1}),\ldots,\ell(x_{n-1}),\ell(y)),$
for all $w\in D_{n}$, is a Lehmer code for type $D_{n}$, for all $n\geqslant 4$.
5.5 A Lehmer code for type $H_{3}$
Let $(H_{3},S)$ be a Coxeter system of type $H_{3}$. We order the set $S$ as in Figure 3. A Lehmer code can be constructed in the following way.
[…]
Figure 3: Coxeter graph of $H_{3}$.
Define three chains $X$, $Y$ and $Z$ in the Bruhat order of $H_{3}$ by setting
$X$ ::=:= $\{e,s_{3},s_{3}s_{2},s_{3}s_{2}s_{3},s_{3}s_{2}s_{3}s_{2},s_{3}s_{2}s_{3}s_{2}s_{1},s_{3}s_{2}s_{3}s_{2}s_{1}s_{2},$
$s_{3}s_{2}s_{3}s_{2}s_{1}s_{2}s_{3},s_{3}s_{2}s_{3}s_{2}s_{1}s_{2}s_{3}s_{2},s_{3}s_{2}s_{3}s_{2}s_{1}s_{2}s_{3}s_{2}s_{3}\},$
$Y$ ::=:= $\{e,s_{2},s_{2}s_{1},s_{3}s_{2}s_{1},s_{2}s_{3}s_{2}s_{1},s_{1}s_{2}s_{3}s_{2}s_{1}\},$
$Z$ ::=:= $\{s_{1},s_{1}s_{2},s_{1}s_{2}s_{1},s_{1}s_{3}s_{2}s_{1},s_{2}s_{1}s_{3}s_{2}s_{1},s_{1}s_{2}s_{1}s_{3}s_{2}s_{1}\}.$
Notice that $X\subseteq{{}^{S\setminus\{s_{3}\}}H_{3}}$. In the following proposition we provide several characterizations of the set $Y\cup Z$. For a Coxeter system $(W,S)$ and a subset $V\subseteq W$, a generalized quotient is the set
$W/V:=\{w\in W:\ell(wv)=\ell(w)+\ell(v),\,\forall\,v\in V\}.$
For definitions and results about generalized quotients in Coxeter groups we refer to [8]. See also [5, Exercise 2.22]. We denote by $[u,v]_{L}$ an interval in the left weak order of $W$.
Proposition 5.9.
Let $x_{0}:=\max X$ and $z_{0}:=\max Z$. Then
$Y\uplus Z=[e,z_{0}]_{L}=(H_{3})_{\{s_{1},s_{2}\}}\{e,s_{3}s_{2}s_{1}\}=H_{3}/{\{x_{0}\}}=H_{3}/X.$
Proof.
Notice that $Y\cap Z=\varnothing$. It can be checked by hand that $Y\cup Z=[e,z_{0}]_{L}=(H_{3})_{\{s_{1},s_{2}\}}\{e,s_{3}s_{2}s_{1}\}$ and that $z_{0}x_{0}=w_{0}$. By [8, Theorem 4.4], we have the equality $H_{3}/{\{x_{0}\}}=[e,z_{0}]_{L}$. Since $X$ is a saturated chain in the right weak order of $H_{3}$, we have that $H_{3}/X=H_{3}/{\{x_{0}\}}$. ∎
The following corollary provides a factorization of a Coxeter group of type $H_{3}$ similar to a parabolic factorization $H_{3}^{J}\cdot(H_{3})_{J}$.
Corollary 5.10.
The Coxeter group $H_{3}$ factorizes as $H_{3}/X\cdot X$.
Proof.
Notice that $\{e,s_{3}s_{2}s_{1}\}X\subseteq{{}^{S\setminus\{s_{3}\}}H_{3}}$. Moreover $|{{}^{S\setminus\{s_{3}\}}H_{3}}|=|H_{3}|/6=20$ and |$|\{e,s_{3}s_{2}s_{1}\}X|=20$, so $\{e,s_{3}s_{2}s_{1}\}X={{}^{S\setminus\{s_{3}\}}H_{3}}$. By Proposition 5.9 we have the equality $Y\cup Z=(H_{3})_{\{s_{1},s_{2}\}}\{e,s_{3}s_{2}s_{1}\}$ and then $(Y\cup Z)X=H_{3}$. ∎
Now we are ready to define a Lehmer code in type $H_{3}$.
Theorem 5.11.
The function $L_{H_{3}}:H_{3}\rightarrow[1]_{0}\times[5]_{0}\times[9]_{0}$ defined by
$L_{H_{3}}(ux)=\left\{\begin{array}[]{ll}(0,\ell(u),\ell(x)),&\hbox{if $u\in Y$;}\\ (1,\ell(u)-1,\ell(x)),&\hbox{if $u\in Z$,}\end{array}\right.$
for all $u\in H_{3}/X$, $x\in X$, is a Lehmer code.
Proof.
By Corollary 5.10 the function $L_{H_{3}}$ is well defined and it is clearly bijective. Since $X$, $Y$ and $Z$ are saturated chains of the Bruhat order, to complete the proof we have only to check that $\ell(y)=\ell(z)-1$ implies $y<z$, for all $y\in Y$, $z\in Z$. This property is easily verified. ∎
In Example 6.8 we show that the $L_{H_{3}}$-unimodal elements, defined in the next section, are in bijection with the palindromic Poincaré polynomials of the lower Bruhat intervals in $(H_{3},\leqslant)$.
Remark 5.12.
We have checked, by using SageMath, that $(H_{3},\leqslant_{R})$ (equivalently $(H_{3},\leqslant_{L})$) does not inject to $[2]\times[6]\times[10]$. This implies that there is no Lehmer code for $H_{3}$ defined in terms of reflections as can be done in types $A_{n}$.
6 $L$-principal and $L$-unimodal elements
Given a Lehmer code $L$ of a finite Coxeter system $(W,S)$, we introduce a partial order $\operatorname{\preccurlyeq_{\mathrm{L}}}$ on $W$ by setting
$u\operatorname{\preccurlyeq_{\mathrm{L}}}v\,\,\Leftrightarrow\,\,L(u)\leqslant L(v),$
for all $u,v\in W$. For $u\operatorname{\preccurlyeq_{\mathrm{L}}}v$, we let $[u,v]_{\mathrm{Lh}}:=\{z\in W:u\operatorname{\preccurlyeq_{\mathrm{L}}}z\operatorname{\preccurlyeq_{\mathrm{L}}}v\}$. We have that $[u,v]_{\mathrm{Lh}}$ is a subposet of $[u,v]$. It is natural to study the elements $w\in W$ for which $[e,w]_{\mathrm{Lh}}=[e,w]$ as sets. For these elements, it is clear that the rank–generating function of $[e,w]$ factorizes as product of $q$-analogs involving the entries of their Lehmer codes. Hence we define a distinguished set of elements of $W$, which we call $L$–principal.
Definition 6.1.
Let $L$ be a Lehmer code for a finite Coxeter system $(W,S)$. An element $w$ is $L$-principal if $[e,w]_{\mathrm{Lh}}=[e,w]$ as sets.
We denote by $\Pr(L)$ the set of $L$-principal elements in $W$. Notice that the poset $(W,\operatorname{\preccurlyeq_{\mathrm{L}}})$ is a lattice. The lattice operations are the ones induced by the lattice $\operatorname{\mathrm{Im}}(L)$, i.e.
$u\curlywedge v=L^{-1}(L(u)\wedge L(v)),\,\,\,u\curlyvee v=L^{-1}(L(u)\vee L(v)),$
for all $u,v\in W$. It turns out that for any Lehmer code also $\Pr(L)$ is a lattice.
Proposition 6.2.
Let $(W,S)$ be a finite Coxeter system with a Lehmer code $L$. Then the set $\Pr(L)\subseteq W$ is a lattice.
Proof.
Let $a_{1},a_{2}\in\Pr(L)$, $a:=a_{1}\curlywedge a_{2}$ and $w\leqslant a$. We have that $a\operatorname{\preccurlyeq_{\mathrm{L}}}a_{i}$, for all $i\in[2]$. Since $a_{i}\in\Pr(L)$, we have that $a\leqslant a_{i}$, for all $i\in[2]$. Hence $w\leqslant a_{i}$, for all $i\in[2]$, i.e. $w\operatorname{\preccurlyeq_{\mathrm{L}}}a_{i}$, for all $i\in[2]$, by the $L$-principality of $a_{i}$. We have proved that $w\operatorname{\preccurlyeq_{\mathrm{L}}}a$ for all $w\leqslant a$, i.e. $a\in\Pr(L)$, and this concludes the proof. ∎
Notice that, by definition of $L$-principality, $(\Pr(L),\leqslant)=(\Pr(L),\operatorname{\preccurlyeq_{\mathrm{L}}})$. In types $A_{n}$ and $B_{n}$, the principal elements for the Lehmer codes $L_{A_{n}}$ and $L_{B_{n}}$ defined in Section 5 are known in literature as Kempf elements.
Example 6.3.
By [18, Propositions 4.3 and 4.9], the elements of $\Pr(L_{A_{n}})$ and $\Pr(L_{B_{n}})$ are exactly the Kempf elements in type $A_{n}$ and $B_{n}$, respectively. It is known that Kempf elements in type $A_{n}$ coincide with $312$–avoiding permutations of $S_{n+1}$, see Section 7 and references therein.
It is clear that for an $L$-principal element $w\in W$ we have that the Poincaré polynomial of $[e,w]$ is
$h_{w}(q)=\prod\limits_{i=1}^{|S|}[L(w)_{i}+1]_{q}.$
In general, for a finite Coxeter system, a more involved formula for the cardinality of $[e,w]$ holds, as we prove in the following theorem.
Theorem 6.4.
Let $L$ be a Lehmer code of $(W,S)$ and $w\in W$. Then
$h_{w}(q)=\sum\limits_{\varnothing\neq X\subseteq\max\{L(v):v\leqslant w\}}(-1)^{|X|+1}\prod\limits_{i=1}^{|S|}[(\wedge_{x\in X}\,x)_{i}+1]_{q}.$
Proof.
For $w\in W$ we have that, as sets,
$[e,w]=\bigcup\limits_{z\in\max\{L(v):v\leqslant w\}}[e,L^{-1}(z)]_{\mathrm{Lh}}.$
Moreover the rank–generating function of the poset $[e,x]_{\mathrm{Lh}}\cap[e,y]_{\mathrm{Lh}}=[e,x\curlywedge y]_{\mathrm{Lh}}$ is equal to the polynomial $\prod_{i=1}^{|S|}[L(x\curlywedge y)_{i}+1]_{q}$, for all $x,y\in W$. By the Inclusion-Exclusion principle we obtain the result. ∎
Example 6.5.
Let $w:=3412\in S_{4}$ and $L:=L_{A_{3}}$. Then $L(w)=(0,2,2)$ and $\max_{\operatorname{\preccurlyeq_{\mathrm{L}}}}[e,w]=\{2413,3214,3412\}$. We have that $2413\curlywedge 3214=2134$, $2413\curlywedge 3412=1423$, $3214\curlywedge 3412=3124$ and $2413\curlywedge 3214\curlywedge 3412=1234$. Therefore
$h_{w}(q)$ $=$ $\prod\limits_{i=1}^{3}[L(2413)_{i}+1]_{q}+\prod\limits_{i=1}^{3}[L(3214)_{i}+1]_{q}$
$+\prod\limits_{i=1}^{3}[L(3412)_{i}+1]_{q}-\prod\limits_{i=1}^{3}[L(2134)_{i}+1]_{q}$
$-\prod\limits_{i=1}^{3}[L(1423)_{i}+1]_{q}-\prod\limits_{i=1}^{3}[L(3124)_{i}+1]_{q}$
$+\prod\limits_{i=1}^{3}[L(1234)_{i}+1]_{q}$
$=$ $2(1+q)(1+q+q^{2})+(1+q+q^{2})^{2}$
$-(1+q)-2(1+q+q^{2})+1$
$=$ $1+3q+5q^{2}+4q^{3}+q^{4}.$
Now we define a subset of $\Pr(L)$ whose elements are involved in the description of palindromic Poincaré polynomials. Given a Coxeter system $(W,S)$ with a Lehmer code $L$, we define the $L$-orbit of $w\in\Pr(L)$ by setting
$O_{w}:=\{L(u):u\in\Pr(L)\}\cap\{(L(w)_{\sigma(1)},\ldots,L(w)_{\sigma(n)}):\sigma\in S_{n}\},$
where $n:=|S|$.
Definition 6.6.
We say that an $L$-principal element $w\in W$ is $L$-unimodal if $L(w)=\min_{\leqslant_{\operatorname{\mathrm{lex}}}}O_{w}$.
We denote by $U(L)$ the set of $L$-unimodal elements in $W$. In Section 7, for the Lehmer code $L_{A_{n}}$ of $S_{n+1}$, we prove that $U(L_{A_{n}})$ is the set of unimodal permutations in $S_{n+1}$.
We observe that the inclusion $U(L)\subseteq\Pr(L)$ implies the poset equality $(U(L),\leqslant)=(U(L),\operatorname{\preccurlyeq_{\mathrm{L}}})$, for every Lehmer code $L$. Given a Coxeter system $(W,S)$ we define
$\operatorname{\mathrm{Pal}}(W,S):=\left\{h_{w}(q):w\in W,\,q^{\ell(w)}h_{w}(q^{-1})=h_{w}(q)\right\}.$
It is clear by definition that
$\{h_{w}(q):w\in U(L)\}=\{h_{w}(q):w\in\Pr(L)\}\subseteq\operatorname{\mathrm{Pal}}(W,S)$ (3)
and $|U(L)|=|\{h_{w}(q):w\in U(L)\}|$, for every Lehmer code $L$ of a finite Coxeter system $(W,S)$. In Section 7 we prove that in type $A_{n}$, for the Lehmer code $L_{A_{n}}$, we have the equalities
$|\operatorname{\mathrm{Pal}}(A_{n})|=|\{h_{w}(q):w\in U(L_{A_{n}})\}|=2^{n},$
for all $n\in\operatorname{\mathbb{N}}$.
In the following example we show that in general the previous equality does not hold.
Example 6.7.
It can be checked by using a computer that $|\operatorname{\mathrm{Pal}}(B_{3})|>|U(L_{B_{3}})|$. The linear order on the generators given in Figure 4 provides a Lehmer code $\tilde{L}_{B_{n}}$ for type $B_{n}$, by a construction similar to the one of Theorem 5.4. In this case we have that $|\operatorname{\mathrm{Pal}}(B_{3})|=|U(\tilde{L}_{B_{3}})|$ but $|\operatorname{\mathrm{Pal}}(B_{4})|>|U(\tilde{L}_{B_{4}})|$. Finally, similar computations show that $|\operatorname{\mathrm{Pal}}(D_{4})|=|U(L_{D_{4}})|$ and $|\operatorname{\mathrm{Pal}}(D_{5})|>|U(L_{D_{5}})|$.
[…]
Figure 4: Coxeter graph of $B_{n}$.
In the next example we classify palindromic Poincaré polynomials of lower Bruhat intervals in type $H_{3}$. This classification relies on the set of unimodal elements with respect to the Lehmer code $L_{H_{3}}$ defined in Section 5.5.
Example 6.8.
Consider the Lehmer code $L_{H_{3}}$ of Theorem 5.11. We have that
$\operatorname{\mathrm{Pal}}(H_{3})=\left\{h_{w}(q):w\in\Pr(L_{H_{3}})\right\}=\left\{h_{w}(q):w\in U(L_{H_{3}})\right\},$
where the first equality can be checked by a SageMath computation and the second equality follows from (3). Figure 5 depicts the poset of all palindromic Poincaré polynomials of lower Bruhat intervals in type $H_{3}$: each element $(x_{1},x_{2},x_{3})$ in this poset corresponds to an $L_{H_{3}}$-unimodal element whose lower Bruhat interval has Poincaré polynomial equal to $[x_{1}+1]_{q}[x_{2}+1]_{q}[x_{3}+1]_{q}$.
${(1,5,9)}$${(1,5,4)}$${(1,4,4)}$${(1,3,4)}$${(1,2,4)}$${(1,1,4)}$${(1,2,3)}$${(0,1,4)}$${(1,1,3)}$${(1,2,2)}$${(0,1,3)}$${(1,1,2)}$${(0,1,2)}$${(1,1,1)}$${(0,1,1)}$${(0,0,1)}$${(0,0,0)}$
Figure 5: Hasse diagram of the Bruhat order of $U(L_{H_{3}})$.
7 Unimodal permutations and Poincaré polynomials of smooth Schubert varieties
In this section we consider the symmetric group $S_{n}$ with its standard Coxeter presentation. For the Lehmer code $L_{A_{n-1}}$, principal elements are exactly Kempf elements in type $A_{n-1}$ and these are the permutations in $S_{n}$ avoiding the pattern $312$. We prove that $U(L_{A_{n-1}})$ is the set of unimodal permutations in $S_{n}$. The main result of this section is the classification of Poincaré polynomials of smooth Schubert varieties in flag manifolds. These are covered by the rank–generating functions for lower Bruhat intervals of unimodal permutations, as we prove in Theorem 7.11.
Let $P(n)$ be the set of partitions of a non–negative integer $n$. To write a partition we adopt French notation. Then an element $\lambda\in P(n)$ is either $(0)$, if $n=0$, or a tuple $(\lambda_{1},\ldots,\lambda_{r})$, where $0<\lambda_{1}\leqslant\lambda_{2}\leqslant\ldots\leqslant\lambda_{r}$ and $\sum_{i=1}^{r}\lambda_{i}=n$, if $n>0$. Given a partition $(\lambda_{1},\ldots,\lambda_{r})\in P(n)$, we define $[\lambda]:=[r]$ and, given $k\in\mathbb{N}$, $m_{k}(\lambda):=|\{i\in[r]:\lambda_{i}=k\}|$. We denote by $\lambda^{*}$ the dual partition (or conjugate partition), i.e. the unique partition determined by the condition $m_{k}(\lambda^{*}):=\lambda_{r-k+1}-\lambda_{r-k}$, for all $k\in[r]$ and where $\lambda_{0}:=0$. The partition $(0)\in P(0)$ is self–dual.
The set of smooth permutations in $S_{n}$ is defined by setting
$\operatorname{\mathrm{Sm}}_{n}:=\{w\in S_{n}:\mbox{$w$ avoids $3412$ and $4231$}\}.$
It is known that $w\in\operatorname{\mathrm{Sm}}_{n}$ if and only if
$h_{w}(q)=\sum\limits_{v\leqslant w}q^{\ell(v)}=\prod\limits_{i\in[E(w)]}[E(w)_{i}+1]_{q},$
where $E:\operatorname{\mathrm{Sm}}_{n}\rightarrow P(\ell(w))$ is the function which assigns to any element $w\in\operatorname{\mathrm{Sm}}_{n}$ its exponents (see [14, Theorem 1.1]). To define $E(w)$, following [14, Remark 2.8], we need some preparation.
For $w\in S_{n}$ define a poset $P_{w}=([n],\leqslant_{w})$ by setting $i<_{w}j$ if and only if $i<j$ and $w^{-1}(j)<w^{-1}(i)$, for all $i,j\in[n]$. The comparability graph of the poset $P_{w}$ is known as a permutation graph. For a finite poset $P$, we define the poset $C(P)$ of saturated chains of $P$, ordered by inclusion. It is a ranked poset; we denote by $\rho$ its rank function and define $\rho_{i}:=|\{c\in C(P):\rho(c)=i\}|$, for all $i\geqslant 0$, and, if $P$ is not trivial, $m_{P}:=\max\{\rho_{i}:i\geqslant 1\}$. For a ranked poset $P$ of cardinality $n$ whose Hasse diagram is a forest, Proposition 7.1 allows to define a partition into distinct parts.
Proposition 7.1.
Let $P$ be a finite poset whose Hasse diagram is a forest. If $0\leqslant h<k\leqslant m_{P}$, then $\rho_{h}>\rho_{k}$.
Proof.
It is sufficient to prove the result for trees. The condition $m_{P}>0$ implies $|P|>1$. We proceed by induction on $|P|$. If $|P|=2$ then $m_{P}=1$ and $\rho_{0}=2>1=\rho_{1}$. Let $|P|>2$. If $|\min P|=|\max P|=1$ then $P$ is a chain and the result is obvious. Assume $|\max P|\neq 1$. Let $C$ be a maximal saturated chain of length $m_{P}$. We claim that there exists a leaf of $P$ not belonging to $C$. Recall that in a tree there are at least two leaves. Suppose first that $P$ has exactly two leaves. In this case it is clear that $C$ cannot have these two leaves as extremal points and the claim follows. If $P$ has more than two leaves, the claim follows immediately. This implies that there exists $x\in P$ such that $m_{P\setminus\{x\}}=m_{P}$. Let $P^{\prime}:=P\setminus\{x\}$. Then $\rho^{\prime}_{j}\in\{\rho_{j}-1,\rho_{j}\}$, for all $j\in[m_{P}]$. We have that $\rho^{\prime}_{h}=\rho_{h}$ implies $\rho^{\prime}_{k}=\rho_{k}$, for all $0\leqslant h<k\leqslant m_{P}$. Hence, if $\rho^{\prime}_{h}=\rho_{h}$, by our inductive hypothesis we have $\rho_{h}=\rho^{\prime}_{h}>\rho^{\prime}_{k}=\rho_{k}$. If $\rho^{\prime}_{h}=\rho_{h}-1$, again by our inductive hypothesis, $\rho_{h}=\rho^{\prime}_{h}+1>\rho^{\prime}_{k}+1\geqslant\rho_{k}$, for all $0\leqslant h<k\leqslant m_{P}$. ∎
Given a smooth permutation $w$, the Hasse diagram of the poset $P_{w}$ is a forest (see [9, Section 3]). For $w\in\operatorname{\mathrm{Sm}}_{n}\setminus\{e\}$, by considering the poset $C(P_{w})$ and Proposition 7.1, we define a partition
$\lambda_{w}:=(\rho_{m_{P}},\ldots,\rho_{1})\in P(\ell(w)).$ (4)
We let $\lambda_{e}:=(0)$. Then the function $E$ is defined by setting $E(w):=\lambda^{*}_{w}$, for all $w\in\operatorname{\mathrm{Sm}}_{n}$.
In (2) we have defined a function $L_{n}:S_{n}\rightarrow\{0\}\times[1]_{0}\times\ldots\times[n-1]_{0}$. By Theorem 5.2, this function satisfies
$L_{n}(w)_{k}=|\{i\in[n]:i<k,\,w^{-1}(i)>w^{-1}(k)\}|,$
for all $k\in[n]$, $w\in S_{n}$. With a slight abuse of language, from now on we say that $L_{n}$ is a Lehmer code for $S_{n}$. In order to describe $L_{n}$-principal elements in Proposition 7.8, we study permutations $w\in S_{n}$ avoiding the pattern $312$. For these permutations we can easily relate the partition $\lambda^{*}_{w}$ to the Lehmer code $L_{n}(w)$.
Proposition 7.2.
Let $w\in S_{n}$ be a $312$–avoiding permutation and $\lambda:=\lambda^{*}_{w}$. Then
$\{\lambda_{h}:h\in[\lambda]\}\cup\{0\}=\{L_{n}(w)_{k}:k\in[n]\}.$
In particular,
$h_{w}(q)=\prod\limits_{k=1}^{n}[L_{n}(w)_{k}+1]_{q}.$
Proof.
If $w\in S_{n}$ avoids $312$ then $w\in\operatorname{\mathrm{Sm}}_{n}$. If $w=e$ the result is trivial. Let $w\neq e$. By definition of the poset $P_{w}$, it is clear that for each inversion $(i,j)$ of $w$ we have a saturated chain $c_{(i,j)}$. Let $k\in[n]$ and define $I_{k}(w):=\{(i,k):i<k,w^{-1}(i)>w^{-1}(k)\}$. For a $312$–avoiding permutation $w$, if $(i,k),(j,k)\in I_{k}(w)$, for some $i<j$, then in the poset of saturated chains $C(P_{w})$ we have that $\rho(c_{(i,k)})>\rho(c_{(j,k)})$. This implies that $|I_{k}(w)|$ is a component of $\lambda_{w}^{*}$, for all $k\in[n]$ such that $I_{k}(w)\neq\varnothing$. Since $|I_{k}(w)|=L_{n}(w)_{k}$ for all $k\in[n]$, the first statement follows. The second one is immediate from the fact that $E(w)=\lambda^{*}_{w}$. ∎
We proceed now to define Fubini words and flag Fubini words. Fubini words are enumerated by Fubini numbers (see [28, Section 3]). Flag Fubini words are enumerate by Catalan numbers and they arise as Lehmer codes of $L_{n}$-principal permutations, as stated in Proposition 7.8.
Definition 7.3.
Let $k\in\operatorname{\mathbb{N}}$. An element $x\in[k-1]_{0}^{k}$ is a Fubini word if $\{x_{1},\ldots,x_{k}\}=\{0,\ldots,\max(x)\}$.
As ordered set partitions, Fubini words provide bases of generalized coinvariant algebras, as treated in [17]. Fubini words have a geometric interest as they realize indexing sets for stratifications of spaces of line configurations, see [20]. Since the dual of a partition into distinct parts is a Fubini word, we obtain the following corollary of Proposition 7.1.
Corollary 7.4.
Let $w\in\operatorname{\mathrm{Sm}}_{n}$. Then the dual partition $\lambda_{w}^{*}$, with a 00 prefix if $\lambda_{w}^{*}\neq(0)$, is a Fubini word.
The next definition introduce the notion of flag Fubini word.
Definition 7.5.
An element $x\in[k-1]_{0}^{k}$ is a flag Fubini word if $(x_{1},\ldots,x_{i})$ is a Fubini word for all $i\in[k]$.
It is immediate to verify the following equivalence.
Proposition 7.6.
Let $x\in[k-1]_{0}^{k}$. Then the following are equivalent:
[…]
    $x$ is a flag Fubini word;
[…]
    $x_{1}=0$ and $x_{i+1}-x_{i}\leqslant 1$, for all $i\in[k-1]$.
We denote by $\hat{\mathrm{F}}_{k}$ the set of flag Fubini words in $[k-1]_{0}^{k}$. Notice that $\hat{\mathrm{F}}_{k}\subseteq\{0\}\times[1]_{0}\times\ldots\times[k-1]_{0}$. Moreover, by Proposition 7.6 and [24, Exercise 6.19(u)], flag Fubini words are enumerated by Catalan numbers, i.e. $|\hat{\mathrm{F}}_{k}|=C_{k}$, for all $k\in\operatorname{\mathbb{N}}$.
Example 7.7.
We have that $\hat{\mathrm{F}}_{3}=\{(0,0,0),(0,1,0),(0,0,1),(0,1,1),(0,1,2)\}$ and
[…]
$(0,1,0,1),(0,0,1,1),(0,1,1,1),(0,1,2,0),(0,0,1,2),$
(0,1,1,2),(0,1,2,1),(0,1,2,2),(0,1,2,3)}.(0,1,1,2),(0,1,2,1),(0,1,2,2),(0,1,2,3)\}.( 0 , 1 , 1 , 2 ) , ( 0 , 1 , 2 , 1 ) , ( 0 , 1 , 2 , 2 ) , ( 0 , 1 , 2 , 3 ) } .
The tuple $(0,2,1,1)$ is a Fubini word which is not flag.
Recall from Section 6 that an element $w\in S_{n}$ is $L$-principal for a Lehmer code $L$ of $S_{n}$ if $[e,w]_{\mathrm{Lh}}=[e,w]$ as sets, and that we denote by $\Pr(L)\subseteq S_{n}$ the set of $L$-principal elements. The content of the following proposition is essentially known in the context of Kempf varieties, see [18]. Since a unifying result is not available in our setting, we think it is useful to state it here with a short proof.
Proposition 7.8.
Let $n\in\operatorname{\mathbb{N}}$ and $w\in S_{n}$. The following are equivalent:
[…]
    $w$ is $L_{n}$-principal;
[…]
    $L_{n}(w)$ is a flag Fubini word;
[…]
    $w$ avoids the pattern $312$.
In particular $|\Pr(L_{n})|=C_{n}$, where $C_{n}$ is the $n$-th Catalan number.
Proof.
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ): by contradiction, assume $L_{n}(w)$ is not flag Fubini. Hence by Proposition 7.6 there exists $i\in[n]$ such that $L_{n}(w)_{i}<L_{n}(w)_{i+1}-1$. Then there exists $a<i$ such that $w^{-1}(i+1)<w^{-1}(a)<w^{-1}(i)$. We have that $s_{i}\in D_{L}(w)$ and then $s_{i}w<w$ in the Bruhat order. But $L_{n}(s_{i}w)_{i}>L_{n}(w)_{i}$ and then $s_{i}w\not\in[e,w]_{\mathrm{Lh}}$, i.e. $[e,w]_{\mathrm{Lh}}\neq[e,w]$, a contradiction.
2)3)2$\Rightarrow$ 3)2 ) ⇒ 3 ): by contradiction, assume that $w$ has the pattern $312$. This means that there exist $b<c<a$ with $a,b,c\in[n]$ such that $w^{-1}(a)<w^{-1}(b)<w^{-1}(c)$. Consider $a,b,c$ such that $(w^{-1}(a),w^{-1}(b),w^{-1}(c))$ is lexicographically minimal. If $a=c+1$, then $[L_{n}(w)]_{a}-[L_{n}(w)]_{a-1}>1$; by Proposition 7.6 this is a contradiction. If $a>c+1$, the minimality of $(w^{-1}(a),w^{-1}(b),w^{-1}(c))$ implies that $w^{-1}(k)>w^{-1}(c)$ for all $c+1\leqslant k\leqslant a-1$. It follows that $[L_{n}(w)]_{a}-[L_{n}(w)]_{a-1}>1$, again a contradiction.
3)1)3$\Rightarrow$ 1)3 ) ⇒ 1 ): by Proposition 7.2 we have that
$|[e,w]|=\prod\limits_{i=1}^{n}(L_{n}(w)_{i}+1)=|[e,w]_{\mathrm{Lh}}|.$
Since $[e,w]_{\mathrm{Lh}}$ is a subposet of $[e,w]$, we obtain that, as sets, $[e,w]_{\mathrm{Lh}}=[e,w]$, i.e. $w$ is $L_{n}$-principal.
The last assertion of our theorem follows from the fact that $312$-avoiding permutations are enumerated by Catalan numbers (see [24, Exercise 6.19(ff)]). ∎
In the following definition, we recall the notion of unimodal permutation. For more details and results involving this class of permutations see for instance [1], [13], [27].
Definition 7.9.
A permutation $w\in S_{n}$ is unimodal if there exists $i\in[n]_{0}$ such that $w(j)<w(j+1)$ for all $j<i$, and $w(j)>w(j+1)$ for all $i<j<n$.
We denote by $U_{n}$ the set of unimodal permutations in $S_{n}$. Since unimodal permutations coincide with permutations avoiding $312$ and $213$, Proposition 7.8 implies that every unimodal permutation is $L_{n}$-principal. In the next proposition we provide a further characterization of unimodal permutations.
Proposition 7.10.
Let $n\in\operatorname{\mathbb{N}}$ and $u\in S_{n}$. The following are equivalent:
[…]
    $u$ is $L_{n}$-unimodal;
[…]
    $L_{n}(u)$ is a sorted Fubini word;
[…]
    $u$ is an unimodal permutation.
Proof.
First we prove that 2)2)2 ) is equivalent to 3)3)3 ). Let $u\in U_{n}$. Since $u$ avoids $312$, by Proposition 7.8 we have that $L_{n}(u)$ is a Fubini word. Moreover
$L_{n}(u)_{i+1}=\left\{\begin{array}[]{ll}L_{n}(u)_{i},&\hbox{if $u^{-1}(i)<u^{-1}(i+1)$;}\\ L_{n}(u)_{i}+1,&\hbox{if $u^{-1}(i)>u^{-1}(i+1)$,}\end{array}\right.$
for all $i\in[n-1]$. Hence $L_{n}(u)$ is a sorted Fubini word.
Let $f:=(0,f_{2},\ldots,f_{n})\in[n-1]^{n}$ be a sorted Fubini word, i.e. $f_{i}\leqslant f_{i+1}$ for all $i\in[n-1]$, and $f_{i}\leqslant i-1$, for all $i\in[n]$. We proceed by induction on $n\geqslant 1$ to prove that $u:=L_{n}^{-1}(f)\in U_{n}$. If $n=1$ there is nothing to prove. Let $n>1$. Then $\tilde{u}:=L^{-1}_{n-1}(0,f_{2},\ldots,f_{n-1})\in U_{n-1}$ by our inductive hypothesis, because $(0,f_{2},\ldots,f_{n-1})$ is still a sorted Fubini word. Since $f$ is a sorted Fubini word, $f_{n}\in\{f_{n-1},f_{n-1}+1\}$. If $f_{n}=f_{n-1}$ we have that $u^{-1}(n)=\tilde{u}^{-1}(n-1)+1$. If $f_{n}=f_{n-1}+1$ we have that $u^{-1}(n)=\tilde{u}^{-1}(n-1)$. In both cases $u\in U_{n}$.
We prove now the equivalence between 1)1)1 ) and 2)2)2 ). Let $u\in U(L_{n})$. Since $u$ is $L_{n}$-principal, then $L_{n}(u)$ is a flag Fubini word by Proposition 7.8. Since $L_{n}(u)=\min_{\leqslant_{\operatorname{\mathrm{lex}}}}O_{u}$ it follows immediately that $L_{n}(u)$ is a sorted Fubini word. Finally, let $u\in S_{n}$ such that $L_{n}(u)$ is a sorted Fubini word; then it is a flag Fubini word. By Proposition 7.8, $u$ is $L_{n}$-principal and $L_{n}(u)$ is the lexicographically minimum in its orbit because $L_{n}(u)$ is sorted. ∎
In the next result we prove that all the possible Poincaré polynomials of smooth permutations in $S_{n}$ arise as Poincaré polynomials of unimodal permutations.
Theorem 7.11.
Let $n\in\operatorname{\mathbb{N}}$. Then
$\{h_{w}(q):w\in\operatorname{\mathrm{Sm}}_{n}\}=\{h_{w}(q):w\in U_{n}\}.$
In particular, |$|\{h_{w}(q):w\in\operatorname{\mathrm{Sm}}_{n}\}|=2^{n-1}$.
Proof.
The inclusion $\{h_{w}(q):w\in U_{n}\}\subseteq\{h_{w}(q):w\in\operatorname{\mathrm{Sm}}_{n}\}$ is trivial because every unimodal permutation avoids the patterns $3412$ and $4231$.
Let $w\in\operatorname{\mathrm{Sm}}_{n}$. Consider the dual partition $\lambda^{*}_{w}$ of the partition $\lambda_{w}$ defined in (4). Define $\widetilde{\lambda}_{w}$ equal to $\lambda^{*}_{w}$ if $w=e$ and to $\lambda^{*}_{w}$, to which we added as a prefix $n-[\lambda_{w}]$ zeroes, if $w\neq e$. By Corollary 7.4, $\widetilde{\lambda}_{w}$ is a sorted Fubini word. By Proposition 7.10, $u:=L^{-1}_{n}(\widetilde{\lambda}_{w})$ is a unimodal permutation. By [14, Remark 2.8] and Proposition 7.2, $h_{w}(q)=h_{u}(q)$. The last assertion is straightforward because $L_{n}$ is a bijection and $|U_{n}|=2^{n-1}$. ∎
Unimodal permutations of length $k$ in $S_{n}$ are in bijection with partitions $(\lambda_{1},\ldots,\lambda_{r})$ of $k$ such that $\lambda_{1}<\ldots<\lambda_{r}\leqslant n-1$, see [1, Theorem 10]. We denote by $\Lambda$ this bijection and then by $\Lambda(u)$ the partition associated to $u\in U_{n}$. We recall how this bijection works. We let $\Lambda(e):=(0)$. Given $u\in U_{n}\setminus\{e\}$, let $j:=u^{-1}(n)$. Then
$\Lambda(u)=(n-u(j+1),\ldots,n-u(n)).$
Conversely, given a partition $\lambda=(\lambda_{1},\ldots,\lambda_{r})$ of $k$, such that $\lambda_{1}<\ldots<\lambda_{r}\leqslant n-1$, consider the set $\{t_{1},\ldots,t_{n-r}\}:=[n]\setminus\{n-\lambda_{1},\ldots,n-\lambda_{r}\}$ and assume that $t_{1}<\ldots<t_{n-r}$. Then
$\Lambda^{-1}(\lambda):=t_{1}\ldots t_{n-r}\,\,n-\lambda_{1}\ldots n-\lambda_{r}.$
The last result of the paper allows to obtain the partition in (4), for unimodal permutations, in a considerably simplified way by using the bijection $\Lambda$.
Proposition 7.12.
Let $u\in U_{n}$. Then $\Lambda(u)=\lambda_{u}$ and
$(\underbrace{0,\ldots,0}\limits_{u(n)},\Lambda(u)^{*}_{1},\ldots,\Lambda(u)^{*}_{s})=L_{n}(u),$
where $s=n-u(n)$ is the number of parts of $\Lambda(u)^{*}$.
Proof.
Let $\Lambda(u)=(\lambda_{1},\ldots,\lambda_{r})$ and $\Lambda(u)^{*}=(\mu_{1},\ldots,\mu_{s})$. We prove that $(0,\ldots,0,\mu_{1},\ldots,\mu_{s})$ is equal to $L_{n}(u)$. If $\ell(u)=0$ we have nothing to show. Let $\ell(u)>0$. It is clear that $L_{n}(u)_{i}=0$ for all $i\leqslant u(n)$ and $[L_{n}(u)]_{u(n)+1}\neq 0$. We have that $s=\lambda_{r}=n-u(n)$ and the number of zeros in $L_{n}(u)$ is $u(n)=n-s$.
Let $0<k\leqslant n-u^{-1}(n)=r$; since $u$ is unimodal, we have that
|$|\{i\in[n]:L_{n}(u)_{i}=k\}|=u(n-k)-u(n-k+1).$
Hence, by defining $\lambda_{0}:=0$, we obtain
$u(n-k)-u(n-k+1)$ $=$ $n-u(n-k+1)-(n-u(n-k))$
$=$ $\lambda_{r-k+1}-\lambda_{r-k}$
and this concludes the proof of the second equality by definition of dual partition.
By [14, Remark 2.8] we have that $\lambda_{u}^{*}$ provides the exponents of $u$. By Proposition 7.2 also the Lehmer code $L_{n}(u)$ provides the exponents of $u$. Hence, by the equality proved above, $\Lambda(u)^{*}=\lambda_{u}^{*}$ and then $\Lambda(u)=\lambda_{u}$. ∎
By using Theorem 7.11 and Proposition 7.12, it is possible to list easily all the possible Poincaré polynomials of smooth permutations in $S_{n}$, as the following example shows.
Example 7.13.
The partitions corresponding to the elements of $U_{4}$ are
$(1,2,3),(2,3),(1,3),(1,2),(3),(2),(1),(0).$
Then the elements of the set $\{h_{w}(q):w\in\operatorname{\mathrm{Sm}}_{4}\}$ are
$[2]_{q}[3]_{q}[4]_{q},\,\,[2]_{q}[3]_{q}[3]_{q},\,\,[2]_{q}[2]_{q}[3]_{q},\,\,[2]_{q}[3]_{q},\,\,[2]_{q}[2]_{q}[2]_{q},\,\,[2]_{q}[2]_{q},\,\,[2]_{q},\,\,[1]_{q}.$
Acknowledgements: The authors would like to thank the organizers of the Workshop “Bruhat order: recent developments and open problems”, held at the University of Bologna on April 2024 where part of our research was carried on.